OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2.

OpenStudy (anonymous):

I know 2 and 3 are the first numbers but what about the rest?

OpenStudy (anonymous):

for f(3)... n=3... then f(3)=f(3-1)+f(3-2) so.. f(3)=f(2)+f(1)

OpenStudy (anonymous):

What is the sequence?

OpenStudy (anonymous):

what is the value of f(3) ?

OpenStudy (anonymous):

How would you find that?

OpenStudy (anonymous):

I don't know the value of f

OpenStudy (anonymous):

put n=3 in f(n)=f(n-1)+f(n-2)

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

Or is it 4?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

with n=3.. then f(n) = f(3)

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

alright?

OpenStudy (anonymous):

Then is it 8?

OpenStudy (anonymous):

wait... the answer is 5.. but did you understand ?

OpenStudy (anonymous):

Yes, kind of. :) I thought about it. :)

OpenStudy (anonymous):

hehe then f(4)= ?

OpenStudy (anonymous):

Is it 8?

OpenStudy (anonymous):

yeah... great

OpenStudy (anonymous):

Ok, thanks sooo much! :) That's all I needed. :)

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

:)